OpenStudy (anonymous):

Find the anti-derivative of x(x^2+1)^-1/2

OpenStudy (anonymous):

\[\frac{ x }{ \sqrt{x^2+1} }\]

OpenStudy (solomonzelman):

set u=x^2+1

OpenStudy (solomonzelman):

I will give you a latex tip. `\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{x}{ \sqrt{x^2+1} }~dx}\)` this I use, to put up: \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{x}{ \sqrt{x^2+1} }~dx}\)

OpenStudy (solomonzelman):

but, really, try \(\large\color{slate}{\displaystyle u=x^2+1}\)

OpenStudy (solomonzelman):

what will you get for du ?

OpenStudy (anonymous):

2x

OpenStudy (solomonzelman):

yes, so \(\large\color{slate}{\displaystyle du=2x~dx}\)

OpenStudy (solomonzelman):

you had: \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{x}{ \sqrt{x^2+1} }~dx}\). Now do you substitution.

OpenStudy (anonymous):

I didn't think you could find the derivative of each individual part?....

OpenStudy (solomonzelman):

you can't

OpenStudy (anonymous):

So you're saying... \[\frac{ x }{ \sqrt{u} }\] ?

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

there are two ways to write this. way 1 \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{x}{\sqrt{x^2+1}}dx}\) \(\large\color{gray}{\displaystyle u=x^2+1}\) \(\large\color{gray}{\displaystyle du=2x~dx~~~~~\Longrightarrow~~~~~dx=\frac{du}{2x}}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{x}{\sqrt{u}}\frac{du}{2x}}\) the x's cancel and you take the 1/2 constant out. \(\large\color{black}{\displaystyle \frac{1}{2}\int\limits_{~}^{~}\frac{1}{\sqrt{u}}du}\)

OpenStudy (solomonzelman):

I keep disconnecting sorry

OpenStudy (anonymous):

All good. I'm still soaking in the last response. Trying to comprehend.

OpenStudy (solomonzelman):

way 2. \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{x}{\sqrt{x^2+1}}~dx}\) \(\large\color{red}{\displaystyle u=x^2+1}\) \(\large\color{slate}{\displaystyle du=2x~dx ~~~~~~~\Rightarrow}\) \(\large\color{green}{\displaystyle \frac{1}{2}du=x~dx }\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{\color{green}{x }}{\sqrt{\color{red}{\displaystyle x^2+1}}}~\color{green}{dx }}\) substitution, \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{\color{green}{\frac{1}{2} }}{\sqrt{\color{red}{\displaystyle u}}}~\color{green}{du }}\) take the constant out, \(\large\color{slate}{\displaystyle\color{green}{\frac{1}{2}}\int\limits_{~}^{~}\frac{\color{green}{1 }}{\sqrt{\color{red}{\displaystyle u}}}~\color{green}{du }}\)

OpenStudy (solomonzelman):

you get the exact same thing, but a substitution is conceptualized a little differently.

OpenStudy (solomonzelman):

I think 1st approach is more standard though, so use that one I would advise.

OpenStudy (anonymous):

Right... so now I would take the anti-derivative of 1/sqrtu in terms of u? The second example is easier to understand.. thank you.

OpenStudy (anonymous):

Why is du = 2xdx?

OpenStudy (solomonzelman):

oh, if you want 2nd approach, then it would also work..... And then the anti-derivative of u is not hard. re-write it as u^(-1/2)

OpenStudy (anonymous):

Nevermind. Got it.

OpenStudy (solomonzelman):

oh, ok.

OpenStudy (anonymous):

Okay... anti-derivative of u^-1/2 would be 2sqrtu

OpenStudy (anonymous):

No. The two would cancel?

OpenStudy (anonymous):

So the final answer is \[\sqrt{x^2+1}\]

OpenStudy (solomonzelman):

yes, the expression is good. But you left out the +c

OpenStudy (anonymous):

Darn. I always do that. Lol

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle \frac{d}{dx}~(x^2+1)^{1/2}+C= \\ ~\\~\\ \frac{1}{2}(x^2+1)^{1/2~~-1~~}\color{blue}{\times~2x}\\ ~\\~\\ \frac{1}{2}(x^2+1)^{-1/2}\color{blue}{\times~2x}\\ ~\\~\\ (x^2+1)^{-1/2}\color{blue}{\times~x}~\\~\\ \displaystyle\frac{x}{\sqrt{x^2+1}}}\)

OpenStudy (solomonzelman):

yes, just checked...

OpenStudy (solomonzelman):

(yes, the blue is the chain rule)

OpenStudy (anonymous):

I did the same. Looks great. Thanks again.

OpenStudy (solomonzelman):

sure:)